OpenStudy (anonymous):

can anyone help me checking my answers in multiplying algebraic expressions?...

OpenStudy (anonymous):

\[1/2(2x^3y^4z^5)(3x^2y^6)\]

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

multiply the given algebraic expression

OpenStudy (anonymous):

Remember when you multiply powers of the same base you add their exponents.. \[(x^a)(x^b) = x^{(a+b)}\]

OpenStudy (anonymous):

So just re-arrange your product so that you're multiplying your bases to eachother, then add their exponents.

OpenStudy (amistre64):

what answer didja get? so we can chk it :)

OpenStudy (anonymous):

i know my ans. is wrong that's why i need your help..

OpenStudy (anonymous):

Ok well did what I said make sense?

OpenStudy (anonymous):

Remember that in a product you can always re-order your factors to make it easier.

OpenStudy (anonymous):

is it (1/2)(2)(3)x^5y^10z^5=3x^5y^10z^5

OpenStudy (anonymous):

So lets start with that: \[\frac{1}{2}(2x^3y^4z^5)(3x^2y^6) \]First we'll get all our constant numbers out in front..\[= (\frac{1}{2}\cdot 2\cdot3)(x^3y^4z^5)(x^2y^6)\] That should simplify nicely.. Now lets have you simplify the constant coefficients and also reorder each of the variables such that they are adjacent to the other power of the same base.

OpenStudy (anonymous):

so, am i going to use vertical method in multiplying \[(x^3y^4z^5)(x^2y^6) \]?

OpenStudy (anonymous):

No, I'd suggest just moving around the factors like we did for the numbers: \[(x^3y^4z^5)(x^2y^6) = (x^3x^2)(y^4y^6)z^5\] Then just remember that when you multiply powers of the same base you add their exponents: \[x^ax^b = x^{a+b}\]

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

I'm heading to bed. I hope that helps. Otherwise we can work more on it tomorrow maybe.

OpenStudy (anonymous):

so it will become like this \[=(x^5)(y^10)z^5\] \[=x^5y^10z^5\]

OpenStudy (anonymous):

that is y raise to 10

OpenStudy (anonymous):

so the final ans. is 3x^5y^10z^5????

OpenStudy (radar):

Pull your answers together in a final answer. The numerical values first then the variables like: (1/2)(3)(2)=3 for a final answer of: \[3x ^{5}y ^{10}z ^{5}\]

OpenStudy (anonymous):

That's right.. Note that when writing complicated exponents (more than 1 character) you have to use the {} to surround the value: y^{10} = \(y^{10}\)